OpenStudy (anonymous):

Under the translation T(-7, 3) the point (1, 6) will become (-6, -3). I said True.

OpenStudy (anonymous):

no, it would be (-6, 9)

OpenStudy (anonymous):

But -7 - 1 = -6 and 3-6 = -3?

OpenStudy (anonymous):

1-7=-6 and 6+3=9

OpenStudy (anonymous):

Oh. Okay, thank you!

OpenStudy (anonymous):

np